# Capture Stage Environments: A Guide to Better Matting

### v1

@@ 5.3 Learning Capacity Student Model
without fine-tuning 17.892 3.209 6.068
with fine-tuning 9.098 2.126 5.126
@@ 5.4 Downstream Application
We demonstrate the improvements of our fine-tuned matting approach on a reconstruction of our capture stage content by optimizing an Instant-NGP model [19] provided by Nerfstudio [23] in its default setting for the bounded model with a maximum grid resolution of $2^{16}$, and optimized for 100 000100000100\,000100 000 iterations. Fig. 7 shows the resulting reconstruction, using the base version of BGMV2 [14] and its fine-tuned counterpart, demonstrating that the fine-tuned model achieves a more faithful reconstruction of the scene.
@@ 7 Conclusion
In this work, we addressed the unique challenges faced by data-driven matting methods in capture-stage environments, and proposed guidelines for a system design to create conditions that best support accurate matting. Alongside technical considerations, we introduced a lightweight knowledge-distillation pipeline that uses sparse scribble annotations to enhance a high-capacity offline model and transfer its accuracy to a real-time implementation, blending offline precision and real-time efficiency. Through quantitative evaluation and integration into a NeRF reconstruction pipeline, we demonstrated both significant improvements in alpha-mask estimation and downstream applications.
Figure 9: The ratio of sparsely annotated data during training significantly affects performance. Training exclusively on scribbled data is not included in the plotted range, resulting in an MSE of $78\cdot 10^{-4}$ and an SAD of $11\cdot 10^{-3}$.

### v2

@@ 5.2 Baseline Methods and Comparisons
MODNet* [7] $258.070$ $28.185$ $11.236$
VideoMatte* [14] $124.566$ $19.31$ $9.955$
BiMatting* [21] $688.757$ $79.614$ $15.941$
BGMV2* (Base) [14] $20.694$ $3.564$ $4.772$
BGM [22] $83.970$ $11.806$ $12.548$
Bg-DAMNet [26] $63.419$ $18.122$ $23.057$
Bg-DAMNet (da) [26] $52.025$ $16.560$ $21.079$
Bg-ViTMatte [30] $7.805$ $5.760$ $4.221$
Bg-ViTMatte (Ours) $4.978$ $1.415$ $3.978$
BGMV2* (D.) (Ours) $8.054$ $2.032$ $3.783$
Table 1: Matting accuracy on capture-stage data. The best results are highlighted in red, the second best in orange. * denotes real-time methods, and indicates models trained on a hybrid dataset. BGMV2* (D.) is the student model trained via knowledge distillation. The top section lists methods without background information, the middle section includes those using background information, and the bottom section shows our proposed methods.
Qualitative Evaluation
To further analyze the impact of the task-specific refinements, we performed a qualitative comparison of both the teacher and student models before and after these refinements. From Fig. 7 we can observe that the teacher model tends to produce false positives in its base configuration, which is mitigated through the refinement. Meanwhile, the student model rather struggles to capture all foreground regions accurately in its base setup. Fine-tuning on the teacher-generated data, however, significantly improves its performance by more precisely outlining foreground elements and reducing errors.
5.3 Ablation Study on Student Model Training
To assess the effectiveness of different training strategies for real-time matting under limited supervision, we conduct an ablation study comparing several training paradigms for the student model (BGMV2). We evaluate how each variant performs in terms of alpha mask accuracy and consistency with the teacher model.
Table 2: Ablation of student model training strategies. Performance is measured on the capture stage validation set.
Training Strategy MSE ($\downarrow$) SAD ($\downarrow$)
Scribble supervision only 89.2 $\cdot 10^{-4}$ 10.4 $\cdot 10^{-3}$
Frozen refinement (2-stage) 27.8 $\cdot 10^{-4}$ 5.75 $\cdot 10^{-3}$
Knowledge distillation 8.05 $\cdot 10^{-4}$ 2.03 $\cdot 10^{-3}$
Training Strategy Comparison
We first fine-tune the base model of BGMV2 [14] directly on our hybrid dataset combining Adobe DIM with the 41 scribble-annotated failure cases. We observe poor performance, especially when the refinement stage is also trained on sparse labels, as shown in Tab. 2. To mitigate the shortcomings of sparse labels, we freeze the refinement module and train only the base network on sparse data, then fine-tune the refinement network on the Adobe dataset. While this improves performance, it still underperforms compared to teacher-supervised training. Finally, we fine-tune the student model using full-resolution alpha masks generated by the refined teacher model on 580 new capture stage images. This method leverages strong pseudo-ground-truth and results in significantly improved metrics.
Figure 8: Qualitative comparisons of novel views from InstantNGP using the original alpha masks from BGMV2 [14] (left) and our improved masks (right). The baseline’s erroneous matting cuts out part of the content and produces unfaithful colors to compensate for the wrong inputs. The refined matting yields a notable quantitative improvement, with PSNR increasing from 30.5 dB to 33.5 dB.
Teacher-Student Alignment Analysis
To assess the learning capacity of the student model, we additionally compare its alpha mask predictions directly against those produced by the refined teacher model on a held-out validation set of 519 capture stage images. This evaluation focuses on alignment quality, rather than absolute accuracy against ground truth, and helps confirm whether the student successfully learns structural cues from the teacher. Table 3 shows a consistent improvement in similarity between student and teacher outputs after distillation. These results indicate that the student effectively approximates the teacher’s outputs, even under real-world capture conditions, validating the success of our distillation-based training pipeline.
MSE ($\downarrow$) SAD ($\downarrow$) Grad ($\downarrow$)
$\cdot 10^{-4}$ $\cdot 10^{-3}$ $\cdot 10^{-5}$
BGMV2 (Base) 17.892 3.209 6.068
BGMV2 (Distill.) 9.098 2.126 5.126
@@ 5.4 Downstream Application
We demonstrate the improvements of our fine-tuned matting approach on a reconstruction of our capture stage content by optimizing an Instant-NGP model [19] provided by Nerfstudio [24] in its default setting for the bounded model with a maximum grid resolution of $2^{16}$, and optimized for 100 000100\,000 iterations. Fig. 8 shows the resulting reconstruction, using the base version of BGMV2 [14] and its fine-tuned counterpart, demonstrating that the fine-tuned model achieves a more faithful reconstruction of the scene.
@@ 7 Conclusion
In this work, we addressed the unique challenges faced by data-driven matting methods in capture-stage environments, and proposed guidelines for designing systems that better support accurate matting. Alongside technical considerations, we introduced a lightweight knowledge-distillation pipeline that uses sparse scribble annotations to enhance a high-capacity offline model and transfer its accuracy to a real-time implementation, blending offline precision and real-time efficiency. Through quantitative evaluation and integration into a NeRF reconstruction pipeline, we demonstrated both significant improvements in alpha-mask estimation and downstream applications.